# DisastIR: A Comprehensive Information Retrieval Benchmark for Disaster Management

### v1

@@ Abstract
Kai Yin1  Xiangjue Dong1  Chengkai Liu1  Lipai Huang1 Yiming Xiao1Zhewei Liu2Ali Mostafavi1James Caverlee1 1Texas A&M University  2University of Toronto {kai_yin,xj.dong,liuchengkai,lipai.huang,yxiao,mostafavi,caverlee}@tamu.edu zwei.liu@utoronto.ca
@@ 3.4 User Query Generation
where $LLM_{info}$ and $LLM_{query}$ are LLMs prompted to generate retrieval information needs statements and the query-passage pair respectively, $P_{IN}$ and $P_{QG}$ are prompts for information needs and query generation, $P_{seed}$ is the domain passage, $q$ is the synthesized user query, and $psg$ is the corresponding relevant passage.
To ensure generated queries align with the core characteristics and objectives of each search intent, we design intent-specific prompts for both stages of query generation. The full prompt templates for each intent are provided in Appendix B.
@@ 3.6 Relevance Labeling
To this end, we design three diverse prompts for each search intent and use a single LLM to generate relevance scores. The prompts, inspired by Thomas et al. (2024); Farzi and Dietz (2024); Rahmani et al. (2025), are: (1) zero-shot direct scoring—a single-pass judgment; (2) chain-of-thought reasoning—a multi-step prompt mimicking human-style reasoning; and (3) multi-dimensional attribute scoring—relevance decomposed into interpretable sub-criteria. For each search intent, relevance is defined to align with its specific objectives, reflecting the varying interpretations of “relevance” across different task types (Dai et al., 2022). Full prompt templates are provided in Appendix C.
@@
inf-retriever-v1 7B XL 74.23 67.82 69.33 68.91 53.10 77.70 68.52 66.90 2.36
SFR-Embedding-Mistral 7B XL 71.76 67.58 70.42 70.62 50.86 73.61 67.47 66.75 1.08
inf-retriever-v1-1.5b 1.5B XL 70.64 64.42 66.11 66.53 53.86 76.19 66.29 65.54 1.13
NV-Embed-v2 7B XL 74.77 69.74 43.18 68.64 58.73 77.01 65.34 64.45 1.36
multilingual-e5-large 560M Large 67.29 64.25 63.75 60.26 51.02 75.06 63.61 62.08 2.41
multilingual-e5-large-instruct 560M Large 68.39 64.90 63.24 67.21 49.38 64.31 62.91 62.06 1.34
e5-small-v2 33M Small 65.87 63.00 60.89 62.04 47.09 74.83 62.29 60.91 2.22
e5-base-v2 109M Medium 65.77 63.07 58.37 62.28 45.54 74.64 61.61 60.23 2.25
e5-large-v2 335M Large 60.15 63.42 56.20 62.29 50.96 74.99 61.34 60.20 1.85
NV-Embed-v1 7B XL 68.36 63.02 56.84 60.04 48.31 67.86 60.74 59.24 2.47
granite-embedding-125m 125M Medium 64.90 61.04 47.14 62.78 48.04 71.94 59.31 58.67 1.06
llmrails-ember-v1 335M Large 64.82 63.34 45.93 61.06 44.56 73.99 58.95 58.35 1.02
arctic-embed-m-v2.0 305M Medium 61.44 62.50 47.80 58.04 42.34 65.28 56.23 54.75 2.64
mxbai-embed-large-v1 335M Large 64.57 62.96 40.58 58.47 40.22 68.81 55.93 55.42 0.92
gte-base-en-v1.5 137M Medium 60.66 56.06 47.06 52.49 39.88 71.27 54.57 53.71 1.57
gte-large-en-v1.5 434M Large 67.70 58.55 40.31 53.09 34.81 67.30 53.63 53.09 1.01
snowflake-arctic-embed-l-v2.0 568M Large 55.34 59.43 38.75 60.42 41.19 63.38 53.09 51.90 2.23
bge-base-en-v1.5 109M Medium 49.67 53.97 47.03 58.47 35.58 64.43 51.52 50.71 1.57
bge-large-en-v1.5 335M Large 57.02 54.72 32.71 55.22 35.19 65.14 50.00 48.99 2.03
gte-Qwen2-7B-instruct 7B XL 62.47 39.12 41.40 29.98 47.28 70.98 48.54 47.07 3.03
bge-small-en-v1.5 33M Small 47.38 45.51 27.82 52.13 27.20 56.86 42.82 42.07 1.75
snowflake-arctic-embed-m-v1.5 109M Medium 25.73 30.56 18.31 48.24 43.04 64.95 38.47 36.25 5.79
snowflake-arctic-embed-s 33M Small 36.05 27.46 18.17 42.52 37.27 66.59 38.01 33.59 11.63
snowflake-arctic-embed-l 335M Large 40.82 30.41 15.32 32.70 34.62 56.82 35.11 31.35 10.72
Linq-Embed-Mistral 7B XL 35.61 31.41 26.55 39.46 27.30 45.69 34.34 33.52 2.38
thenlper-gte-base 109M Medium 9.22 5.35 38.54 60.80 42.85 46.64 33.90 31.78 6.24
snowflake-arctic-embed-m 109M Medium 33.35 14.25 8.64 35.30 38.93 56.88 31.23 28.83 7.67
all-mpnet-base-v2 109M Medium 15.06 9.77 16.17 46.30 27.38 37.23 25.32 25.05 1.08
e5-mistral-7b-instruct 7B XL 21.65 19.51 19.48 31.04 20.39 39.57 25.27 24.05 4.82
gte-Qwen2-1.5B-instruct 1.5B XL 13.98 22.21 19.61 23.90 18.00 31.20 21.48 21.27 0.98
@@
Table 3: Performances of 30 evaluated IR models in DisastIR. Models are ranked by their overall performance under exact search (highest to lowest) in DisastIR. “Size Bin” indicates its model parameter size bin category (small, medium, large, and extra large as defined in Appendix G). “TW” represents Twitter. Overall performance across all queries under exact and ANN search are in “Ex. Avg” and “ANN Avg” columns. “Drop” shows the percentage decrease from exact to ANN average scores. Bold indicates the highest value, and underline indicates the second-highest.
@@ 4.1 Query and Passage Characteristics
As shown in Table 1, Twitter-related search tasks tend to have a higher average number of query-passage pairs per query. The candidate pool for each query is built by merging the top 10 passages retrieved by 30 different models. This larger pool in Twitter tasks suggests greater divergence in model outputs, indicating lower agreement among retrieval models when ranking passages in social media contexts within disaster management scenarios. Additional analyses of labeled query-passage pairs are provided in Appendix D.
@@
Since relevance scores in DisastIR are judged by LLM, it is vital to evaluate their consistency with human annotations. Thus, we construct the LVHL dataset (LLM-based Vs. Human Labeling) by sampling disaster management-related query-passage pairs with human-labeled relevance scores from several open-source datasets. MS MARCO (Bajaj et al., 2016) and TriviaQA (Joshi et al., 2017) are for QA, ALLNLI (sentence-transformers, 2021) and XNLI (Conneau et al., 2018) for NLI, Climate-Fever (Diggelmann et al., 2021) for FC, and STSB (Cer et al., 2017) for STS. Appendix E provides details on the construction of LVHL.
@@
To evaluate whether LLM-generated queries can serve as a reliable alternative to human-authored ones for retrieval benchmarking, we construct LVHQ (LLM Vs. Human-generated Query), a comparison set spanning all 48 retrieval tasks. For each task, both an LLM-generated and a human-written query are created based on the same domain passage. All query-passage pairs are annotated using the same method as in DisastIR. Appendix F provides full details on the construction of LVHQ.
@@ 5.1 Models
DisastIR is adopted to comprehensively evaluate open-source IR models and support the selection of suitable IR models for real-world disaster management applications. Models are chosen based on three criteria: (1) strong performance on the MTEB retrieval benchmark; (2) inclusion in widely adopted embedding families such as BGE (Chen et al., 2024; Xiao et al., 2024), E5 (Wang et al., 2022, 2023), Snowflake Arctic (Merrick, 2024), and GTE (Li et al., 2023; Zhang et al., 2024a), which are commonly used as baselines and in downstream IR tasks (Sun et al., 2024; Xu et al., 2024; Lee et al., 2024b, a; Cao, 2025; Park et al., 2025); and (3) coverage of different types of search tasks adapted from Wang et al. (2024b).
We select 30 models with parameter sizes ranging from 33 million to 7 billion. Detailed descriptions of these models and their implementations are provided in Appendix G.
@@ 5.2 Evaluation
We evaluate model performance under two retrieval settings – exact and ANN – using Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) as the primary metric, consistent with prior works.
@@
For large-scale corpora, brute-force retrieval is computationally infeasible. A common solution is a multi-stage architecture, where an ANN search retrieves a candidate set of passages, which are then re-ranked for final output (Tu et al., 2020; Macdonald and Tonellotto, 2021). To reflect real-world large-scale disaster information retrieval scenarios, we also evaluate model performance during the candidate generation stage using ANN search. We adopt Usearch (Vardanian, 2023), a high-performance, memory-efficient library based on the HNSW (hierarchical navigable small world) algorithm (Malkov and Yashunin, 2018), to retrieve top $k$ passages per query using precomputed embeddings. For fair comparison, $k$ is set to match the value used in exact search.
@@ 6.1 Overall Performance
Table 3 summarizes the overall performance of all 30 evaluated models across all queries in DisastIR, with detailed results for each search task provided in Appendix H. The inf-retriever-v1 model ranks highest in both exact and ANN search settings, followed closely by SFR-Embedding-Mistral (1.52% and 0.22% lower, respectively). Among non-XL models, multilingual-e5-large performs best, reaching 92.83% and 92.79% of the top model’s performance. Notably, the lightweight e5-small-v2 model (33M parameters) achieves 90.91% and 91.05% of the top model’s performance, despite being 212 times smaller in size.
The Snowflake-arctic-embed-s model shows the largest performance drop (11.63%) under ANN search compared to exact search (Table 3). Most models exhibit drops within 5%; only five exceeded this margin, four of which belong to snowflake family, indicating strong robustness when switching from exact to ANN search in DisastIR. Among the top-performing models analyzed above, SFR-Embedding-Mistral has the smallest drop (1.08%), while multilingual-e5-large has the largest (2.41%). All subsequent analyses are based on exact search; analyses under ANN search can be conducted similarly.
Figure 5: Best-performing models in each search task.
@@ 6.2 Performance across all 48 Tasks
Figure 3 presents the performance distribution of all evaluated models across all 48 search tasks. While inf-retriever-v1 achieves the highest average NDCG@10, its median performance is lower than that of SFR-Embedding-Mistral, and it exhibits greater variability across tasks, as reflected by a larger interquartile range (IQR). This suggests that inf-retriever-v1 is less stable across diverse search tasks in DisastIR.
As shown in Figure 5, no single model consistently outperforms others across all 48 tasks. Instead, top performance is distributed among three models: inf-retriever-v1, SFR-Embedding-Mistral, and NV-Embed-v2. inf-retriever-v1 achieves the best results in only nine tasks, primarily in STS. This highlights the complexity and diversity of disaster management-related retrieval tasks and reinforces the need for domain-specific IR models in real-world disaster management scenarios. Appendix I provides additional analyses of model performance across 48 tasks.
@@ 6.3 Comparison with General Domain
Figure 4 compares model rankings in DisastIR and MTEB. Ranking value of each model is based on overall performance in DisastIR and official retrieval scores from the MTEB English leaderboard. The Spearman correlation between the two rankings is 0.225 ($p=0.251$), indicating no significant correlation. This suggests that strong performance on general-domain benchmarks does not guarantee effectiveness in disaster management-related retrieval. For example, models like Linq-Embed-Mistral and snowflake-arctic-embed-l perform well in MTEB but poorly in DisastIR, while models from the E5 family show the opposite trend.
Although inf-retriever-v1 ranks highest in both DisastIR and MTEB, detailed analysis in Section 6.2 show that it achieves top performance in fewer than 20% of DisastIR tasks, mostly those related to STS. Furthermore, when computational resources are limited and large models are impractical to serve, relying solely on MTEB rankings for model selection, such as choosing snowflake-arctic-embed-l, may fail to retrieve critical or relevant content. These discrepancies underscore the necessity of a domain-specific benchmark like DisastIR to guide retrieval model selection across different disaster management-related search tasks.

### v2

@@ Abstract
Kai Yin1   Xiangjue Dong1,*   Chengkai Liu1,*   Lipai Huang1 Yiming Xiao1Zhewei Liu2Ali Mostafavi1James Caverlee1 1Texas A&M University  2University of Toronto {kai_yin,lipai.huang,yxiao,mostafavi,caverlee}@tamu.edu zwei.liu@utoronto.ca *Corresponding Authors: xj.dong@tamu.edu, liuchengkai@tamu.edu
@@ 3.4 User Query Generation
where $LLM_{info}$ and $LLM_{query}$ are LLMs prompted to generate retrieval information needs statements and the query-passage pair respectively, $P_{IN}$ and $P_{QG}$ are prompts for information needs and query generation, $P_{seed}$ is the domain passage, $q$ is the synthesized user query, and psgpsg is the corresponding relevant passage.
To ensure generated queries align with the core characteristics and objectives of each search intent, we design intent-specific prompts for both stages of query generation. The full prompt templates for each intent are provided in Appendix C.
@@ 3.6 Relevance Labeling
To this end, we design three diverse prompts for each search intent and use a single LLM to generate relevance scores. The prompts, inspired by Thomas et al. (2024); Farzi and Dietz (2024); Rahmani et al. (2025), are: (1) zero-shot direct scoring—a single-pass judgment; (2) chain-of-thought reasoning—a multi-step prompt mimicking human-style reasoning; and (3) multi-dimensional attribute scoring—relevance decomposed into interpretable sub-criteria. For each search intent, relevance is defined to align with its specific objectives, reflecting the varying interpretations of “relevance” across different task types (Dai et al., 2022). Full prompt templates are provided in Appendix D.
@@
Linq-Embed-Mistral 7B XL 74.40 70.50 64.22 70.77 52.56 71.35 67.30 66.98 0.48
SFR-Embedding-Mistral 7B XL 71.50 67.34 69.62 70.39 51.08 72.71 66.71 66.39 0.48
inf-retriever-v1 7B XL 72.84 66.92 66.37 65.76 52.02 76.00 66.65 65.98 1.01
inf-retriever-v1-1.5b 1.5B XL 69.47 64.40 63.08 65.49 54.14 73.96 65.09 64.85 0.37
NV-Embed-v2 7B XL 74.55 69.51 42.55 68.39 58.39 76.13 64.92 64.57 0.54
gte-Qwen2-1.5B-instruct 1.5B XL 69.96 59.21 65.21 62.84 55.73 73.61 64.43 64.24 0.29
multilingual-e5-large 560M Large 67.08 64.08 62.99 60.06 51.20 74.14 63.26 62.79 0.74
e5-mistral-7b-instruct 7B XL 65.65 65.16 63.42 67.94 47.68 66.39 62.71 61.99 1.15
multilingual-e5-large-instruct 560M Large 68.14 64.72 62.46 66.96 48.75 63.53 62.43 62.01 0.67
e5-small-v2 33M Small 65.66 62.84 60.10 61.78 47.12 73.93 61.90 61.48 0.68
e5-base-v2 109M Medium 65.54 62.91 57.76 62.11 45.52 73.73 61.26 60.72 0.88
e5-large-v2 335M Large 60.03 63.24 55.48 62.03 50.96 74.09 60.97 60.45 0.85
NV-Embed-v1 7B XL 68.14 62.87 56.13 59.85 48.25 67.11 60.39 59.60 1.31
granite-embedding-125m 125M Medium 64.63 60.85 46.55 62.56 48.11 71.06 58.96 58.60 0.61
gte-Qwen2-7B-instruct 7B XL 70.30 47.65 63.24 31.87 53.88 74.86 56.97 55.99 1.72
snowflake-arctic-embed-m-v2.0 305M Medium 61.28 62.31 47.20 57.84 42.43 64.56 55.94 55.15 1.41
mxbai-embed-large-v1 335M Large 64.37 62.79 40.07 58.30 40.26 67.96 55.62 55.25 0.67
gte-base-en-v1.5 137M Medium 60.46 55.85 46.44 52.34 39.85 70.41 54.22 53.93 0.53
bge-base-en-v1.5 109M Medium 51.65 52.89 46.78 60.13 41.41 68.56 53.57 53.13 0.82
gte-large-en-v1.5 434M Large 67.46 58.37 39.71 52.90 34.79 66.51 53.29 53.21 0.15
snowflake-arctic-embed-l-v2.0 568M Large 55.20 59.29 38.26 60.23 41.23 62.64 52.81 52.32 0.93
bge-large-en-v1.5 335M Large 56.88 54.56 32.32 55.03 35.25 64.43 49.74 49.04 1.41
bge-small-en-v1.5 33M Small 56.87 51.24 25.19 55.30 32.95 64.46 47.67 47.00 1.41
snowflake-arctic-embed-s 33M Small 38.69 28.82 21.43 47.30 40.02 66.95 40.54 38.15 5.90
snowflake-arctic-embed-m-v1.5 109M Medium 25.66 30.43 18.09 48.10 42.98 64.20 38.24 36.85 3.63
snowflake-arctic-embed-l 335M Large 40.73 30.33 15.11 32.60 34.44 56.11 34.89 32.17 7.80
thenlper-gte-base 109M Medium 9.16 5.34 38.06 60.58 42.80 45.99 33.66 32.22 4.28
snowflake-arctic-embed-m 109M Medium 33.26 14.22 8.62 35.16 38.75 56.21 31.02 29.42 5.16
snowflake-arctic-embed-m-long 137M Medium 21.43 10.84 19.49 36.20 41.90 55.00 30.81 29.30 4.90
thenlper-gte-small 33M Small 18.20 9.08 11.04 49.81 37.71 55.47 30.22 29.43 2.61
@@
Table 3: Performances of 30 evaluated IR models in DisastIR. Models are ranked by their overall performance under exact search (highest to lowest) in DisastIR. “Size Bin” indicates its model parameter size bin category (small, medium, large, and extra large as defined in Appendix H). “TW” represents Twitter. Overall performance across all queries under exact and ANN search are in “Ex. Avg” and “ANN Avg” columns. “Drop” shows the percentage decrease from exact to ANN average scores. Bold indicates the highest value, and underline indicates the second-highest. E5-small, base, large-v2, and granite-embedding use knowledge distillation during fine-tuning, which involves additional training signals. Performances across different event types are shown in Table 4.
@@ 4.1 Query and Passage Characteristics
The corpus comprises both original and synthetic passages, with synthetic passages making up only 6.8 % of the total. They come from two sources: 8,000 passages generated from page content ( 3.3 %) and 8,464 document-level summaries serving as proxies for full documents ( 3.5 %). Synthetic passages are introduced to enhance diversity and support varied search intents. Original passages, extracted from PDFs, are formal in style and, when chunked, often miss the broader document context. Some search scenarios, such as Twitter retrieval, require informal text featuring emojis, hashtags, or colloquial expressions, which are largely absent in PDF text. Other scenarios, such as QAdoc, demand whole-document understanding, where LLM-generated summaries provide an effective substitute since full texts typically exceed the input limits of IR models. Evaluation of LLM-generated passages is validated in Appendix B.2.
@@
As shown in Table 1, Twitter-related search tasks tend to have a higher average number of query-passage pairs per query. The candidate pool for each query is built by merging the top 10 passages retrieved by 30 different models. This larger pool in Twitter tasks suggests greater divergence in model outputs, indicating lower agreement among retrieval models when ranking passages in social media contexts within disaster management scenarios. Additional analyses of labeled query-passage pairs are provided in Appendix E.
@@
Since relevance scores in DisastIR are judged by LLM, it is vital to evaluate their consistency with human annotations. Thus, we construct the LVHL dataset (LLM-based Vs. Human Labeling) by sampling disaster management-related query-passage pairs with human-labeled relevance scores from several open-source datasets. MS MARCO (Bajaj et al., 2016) and TriviaQA (Joshi et al., 2017) are for QA, ALLNLI (sentence-transformers, 2021) and XNLI (Conneau et al., 2018) for NLI, Climate-Fever (Diggelmann et al., 2021) for FC, and STSB (Cer et al., 2017) for STS. Appendix F provides details on the construction of LVHL.
@@
We further conduct a controlled in-domain annotation experiment. Specifically, we sample 96 query–passage pairs from the DisastIR labeling pool, covering all 48 intent–event type combinations, and ask three PhD students specializing in disaster management to independently annotate them using the same multi-level relevance scale as in our LLM pipeline. Annotation guidelines are adapted from the chain-of-thought reasoning procedure employed for LLM labeling.
Inter-annotator agreement was substantial, with Fleiss’ kappa reaching 0.777. Comparing LLM-generated scores with human annotations yielded Cohen’s kappa values of 0.681–0.734 across individual annotators, 0.690 when averaged, and 0.803/0.715 under majority vote (minimum/maximum tie-breaking), again demonstrating substantial agreement (Table 6). These findings confirm that LLM-based labeling is not only consistent with external human judgments (as shown in LVHL) but also robustly aligned with expert assessments within the disaster management domain.
@@
To evaluate whether LLM-generated queries can serve as a reliable alternative to human-authored ones for retrieval benchmarking, we construct LVHQ (LLM Vs. Human-generated Query), a comparison set spanning all 48 retrieval tasks. For each task, both an LLM-generated and a human-written query are created based on the same domain passage. All query-passage pairs are annotated using the same method as in DisastIR. Appendix G provides full details on the construction of LVHQ.
@@ 5.1 Models
DisastIR is adopted to comprehensively evaluate open-source IR models and support the selection of suitable IR models for real-world disaster management applications. Models are chosen based on two criteria: (1) strong performance on the MTEB retrieval benchmark; and (2) inclusion in widely adopted embedding model families such as BGE (Chen et al., 2024; Xiao et al., 2024), E5 (Wang et al., 2022, 2023), Snowflake Arctic (Merrick, 2024), and GTE (Li et al., 2023; Zhang et al., 2024a), which are commonly used as baselines and in downstream IR tasks (Sun et al., 2024; Xu et al., 2024; Lee et al., 2024b, a; Cao, 2025; Park et al., 2025).
We select 30 models with parameter sizes ranging from 33 million to 7 billion. Detailed descriptions of these models and their implementations are provided in Appendix H.
@@ 5.2 Evaluation
We evaluate model performance under two retrieval settings, exact and ANN, using Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) as the primary metric, consistent with prior works.
@@
For large-scale corpora, brute-force retrieval is computationally infeasible. A common solution is a multi-stage architecture, where an ANN search retrieves a candidate set of passages, which are then re-ranked for final output (Tu et al., 2020; Macdonald and Tonellotto, 2021). To reflect real-world large-scale disaster information retrieval scenarios, we also evaluate model performance during the candidate generation stage using ANN search. We adopt the HNSW (hierarchical navigable small world) algorithm (Malkov and Yashunin, 2018), to retrieve top $k$ passages per query using precomputed embeddings. For fair comparison, $k$ is set to match the value used in exact search.
Figure 4: Best-performing models in each search task.
@@ 6.1 Overall Performance
Table 3 summarizes the overall performance of all 30 evaluated models across all queries in DisastIR, with detailed results for each search task provided in Appendix I. The Linq-Embed-Mistral model achieves the best performance under exact and ANN search settings, followed closely by SFR-Embedding-Mistral (0.877% and 0.881% lower, respectively).
Among all non-XL models, multilingual-e5-large performs best, reaching 94.0% and 93.00% of the top model’s performance. Notably, the lightweight e5-small-v2 model (33M parameters) achieves 91.98% and 91.17% of the top model’s performance, despite being 212 times smaller in size. The E5-V2 series (Wang et al., 2022) and Granite-Embedding (Awasthy et al., 2025) leverage knowledge distillation during fine-tuning, achieving performance that surpasses all models of the same scale and even outperforms many substantially larger models. This highlights the effectiveness of knowledge distillation in enhancing the performance of smaller models.
The Snowflake-arctic-embed-l model shows the largest performance drop (7.80%) under ANN search compared to exact search (Table 3). Most models exhibit drops within 2%; only five exceeded this margin, four of which belong to snowflake family, indicating strong robustness when switching from exact to ANN search in DisastIR. All subsequent analyses are based on exact search; analyses under ANN search can be conducted similarly.
@@ 6.2 Performance across all 48 Tasks
Figure 3 presents the performance distribution of all evaluated models across all 48 search tasks. All top-5 performance models show great variability across tasks, as reflected by the large interquartile range (IQR). This highlights the limited cross-task robustness of current general domain retrieval models and underscores the need to design methods that enhance cross-task consistency, rather than optimizing solely for higher average performance.
As shown in Figure 4, no single model consistently outperforms others across all 48 tasks. Instead, top performance is distributed among four models: Linq-Embed-Mistral, inf-retriever-v1, SFR-Embedding-Mistral, and NV-Embed-v2. This highlights the complexity and diversity of disaster management-related retrieval tasks and reinforces the need for domain-specific IR models in real-world disaster management scenarios. Appendix J provides additional analyses of model performance across 48 tasks.
Among these top-performing models, only NV-Embed-v2 is accompanied by a public technical report (Lee et al., 2024a), enabling a closer look at its varying performance across intents. Its training data emphasizes fact-checking, NLI, and QA but excludes Twitter, explaining strong results on QA and NLI and weaker performance on Twitter. Template usage further matters: NV-Embed-v2 benefits from alignment between training and inference for QA, QAdoc, FC, and NLI, but lacks such alignment for Twitter. Architecturally, unlike models relying on the final <EOS> token, NV-Embed-v2 employs a latent attention layer, which helps it achieve the highest average NDCG@10 (69.39) among top models when Twitter is excluded.
@@ 6.3 Comparison with General Domain
Figure 5 compares model rankings in DisastIR and MTEB. Ranking value of each model is based on overall performance in DisastIR and official retrieval scores from the MTEB English leaderboard. The Spearman correlation between the two rankings is 0.252 ($p=0.188$), indicating no significant correlation. This suggests that strong performance on general-domain benchmarks does not guarantee effectiveness in disaster management-related retrieval. For example, models in snowflake family perform well in MTEB but poorly in DisastIR, while models from the E5 family show the opposite trend.
Furthermore, when computational resources are limited and large models are impractical to serve, relying solely on MTEB rankings for model selection, such as choosing snowflake-arctic-embed-l, may fail to retrieve critical or relevant content. These discrepancies underscore the necessity of a domain-specific benchmark like DisastIR to guide retrieval model selection across different disaster management-related search tasks.